OpenStudy (anonymous):

Convert -450° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

Can someone please help me with this problem

OpenStudy (anonymous):

welcome to open study please read the code of conduct http://openstudy.com/code-of-conduct

OpenStudy (jdoe0001):

\(\bf radians = \cfrac{degrees \times \pi}{180}\)

OpenStudy (anonymous):

That gives me -7.85? my answer which is - 5pie/2 correct?

OpenStudy (anonymous):

my answer choices are -5pi/2 -5pi/4 5pi/2 -5pi

OpenStudy (anonymous):

im i correct? @jdoe0001

OpenStudy (jdoe0001):

well, yes, -7.85 RADIANS, is correct as far as your choices, your choices are in \(\large \pi \) terms, so use the \(\large \pi\) version without multiplying, just simplifying the fraction for that

OpenStudy (anonymous):

Im confused, what is the correct answer then?

OpenStudy (anonymous):

is the answer -5pi/2? because that equals -7.85

OpenStudy (jdoe0001):

well, you were asked to multiply by \(\boldsymbol{\pi}\) in the calculator and give a radian unit, so -7.85 radians is correct but your choices are in \(\pi\) terms, meaning you are not supposed to multiply by \(\pi\) to match up any of the choices

OpenStudy (jdoe0001):

so to match up any of those choices, you'd need to simplify \(\bf \cfrac{-450 \times \pi}{180}\)

OpenStudy (jdoe0001):

or \(\boldsymbol{\cfrac{-450 \pi}{180}}\)